OpenStudy (anonymous):

Will give medal! Evaluate the function for x = –3. f(x) = 4x – 5

OpenStudy (ivettef365):

just substitute the x for -3, so you have f(-3) = 4(-3) - 5 = -12 -5 = -17

OpenStudy (anonymous):

Simply plug in (-3) for x and solve :)